6-3: The Determinants of Market Interest Rates Expected Interest Rate The real risk-free rate is 2.1%. Inflation is expected to be 2.35% this year, 4.45% next year, and then 2.75% thereafter. The maturity risk premium is estimated to be 0.05(t- 1)%, where t-number of years to maturity. What is the yield on a 7-year Treasury note? Round your answer to two decimal places.
Answer #1

Real Risk-free Rate = 2.10%

Inflation Premium = Average of Inflation over 7 years
Inflation Premium = (2.35% + 4.45% + 5 * 2.75%) / 7
Inflation Premium = 20.55% / 7
Inflation Premium = 2.94%

Maturity Risk Premium = 0.05 * (t - 1)%, here t = 7
Maturity Risk Premium = 0.05 * (7 - 1)%
Maturity Risk Premium = 0.30%

Yield on 7-year Treasury Note = Real Risk-free Rate + Inflation Premium + Maturity Risk Premium + Default Risk Premium
Yield on 7-year Treasury Note = 2.10% + 2.94% + 0.30% + 0.00%
Yield on 7-year Treasury Note = 5.34%